# Non-Hermitian entanglement dip from scaling-induced exceptional criticality

### v1

@@ Abstract
Introduction.– Entanglement scaling behavior provides an established diagnostic for critical phase transitions. Deeply rooted in boundary conformal field theory [1, 2, 3, 4, 5, 6, 7, 8, 9], it universally associates logarithmic entanglement entropy (EE) scaling $S\sim\log L$ with 1D criticality [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. This seminal scaling property has also seen generalizations to higher-dimensional and topological systems, where midgap modes and nodal structures contribute distinctive signatures to the entanglement scaling [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], some with measurable prospects [39, 40, 41, 42, 43, 44, 45, 46].
Of late, established results in critical entanglement scaling have been challenged in non-Hermitian contexts [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In so-called exceptional points (EPs) [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], geometric defectiveness blurs the distinction between occupied and unoccupied bands and probability non-conservation leads to fermionic occupancies effectively greater than one, giving rise to enigmatically negative free-fermion EE through the 2-point function [60, 61, 62, 63, 64, 65, 66]. This is further exacerbated by the non-Hermitian skin effect (NHSE) [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], which produces macroscopically many highly defective eigenstates indexed by complex momenta [51, 95, 96, 97, 98, 99, 100, 101, 102, 103, 73, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127]. Yet, despite these complications, all known non-Hermitian free-fermion critical systems still exhibit the celebrated $S\sim\log L$ scaling.
@@
where $|k\rangle$ is the quasi-momentum basis state and $|\psi_{k,\mu}^{R}\rangle,\langle\psi^{L}_{k,\mu}|$ are respectively the right and left eigenbands [128] of $\mathcal{H}$ with band index $\mu$. We stipulate that fermions occupy an arbitrary set of bands $\mu\in occ.$, such that state occupancy is described by the occupied band projector $P=\sum_{k,\mu\in occ.}|\psi_{k,\mu}^{R}\rangle\langle\psi^{L}_{k,\mu}|=\sum_{k}P_{k}|k\rangle\langle k|$, where $P_{k}=\sum_{\mu\in occ.}\psi_{k,\mu}^{R}[\psi^{L}_{k,\mu}]^{*}$. For an entanglement cut where the region $[x_{L},x_{R}]$ is to be truncated, the biorthogonal  [^1] We used the biorthogonal $\bar{P}$ such as to directly connect to the band occupancy interpretation; some other works [12, 14, 16, 17, 19, 20, 21, 23, 24] used the right eigenstate for both the bra and the ket, with very different results. free-fermion EE is given by [23, 19]
@@
Based on Eq. LABEL:barP, we first review why the EE scales like $\log L$ in Hermitian critical points, as illustrated in Fig. 1(a). Consider a critical point $K_{0}$ (red) where the gap closes but the eigenspace remains full-rank i.e. non-defective. Due to the possible mixing of the occupied and unoccupied eigenstates at gap closure, the occupied band projector $P_{k}$ generically become discontinuous i.e. singular at the critical point. However, due to the discrete momentum spacing (black dots) which scales like $2\pi/L$ (cyan arrows) in a finite lattice, the momentum sum in Eq. LABEL:barP only samples $P_{k}$ down to a distance $\Delta k=k-K_{0}$ bounded below by $\Delta K_{\text{min}}\sim\pi/L$ (brown curve) from the critical point i.e. $\bar{P}_{xx^{\prime}}\sim\int_{\pi/L}\text{e}^{i\Delta k(x^{\prime}-x)}P_{\Delta k}\,d(\Delta k)$ with a UV cutoff $\pi/L$. Since analytic singularities lead to power-law decay in their Fourier coefficients [130, 131]  [^2] In general, the exponent of the power-law decay of the Fourier coefficients is directly dependent on the singularity branching order [130, 149]., we expect $P_{xx^{\prime}}$, which are also the 2-point functions, to become long-ranged and cause substantial entanglement across the cut. Mathematically, this entanglement can be shown to diverge logarithmically with the inverse of the UV cutoff [133, 134], thereby yielding $S\sim\log L$ entanglement scaling  [^3] It can be proven that the $\sim 1/L$ cutoff away from a branching singularity in the symbol $P_{k}$ causes the eigenvalues of $\bar{P}$ to deviate from 00 and $1$ with $\log L$ spacing [133]..
@@
Such $\log L$ entanglement scaling persists even if the critical point is geometrically defective (with completely non-orthogonal eigenstates), as is possible in non-Hermitian systems, as illustrated in Fig. 1(b). For a particular band, the overlap between its left and right eigenstates $\langle\psi^{L}_{k}|$, $|\psi^{R}_{k}\rangle$ can be quantified through the phase rigidity [136, 137, 138, 139] $r(k)=\frac{\langle\psi^{L}_{k}|\psi^{R}_{k}\rangle}{\sqrt{\langle\psi^{L}_{k}|\psi^{L}_{k}\rangle\langle\psi^{R}_{k}|\psi^{R}_{k}\rangle}},$ which is always unity in Hermitian systems, but ranging between $0\leq r(k)\leq 1$ in non-Hermitian systems. In particular, at an exceptional critical point (EP) $K_{0}$ where $r(K_{0})=0$ [137] (dark purple), the occupied and unoccupied eigenstates coalesce and the occupied band projector $P_{k}$ is not just discontinuous, but in fact divergent. Specifically, one or more matrix elements of $P_{k}$ can be shown to diverge with $\Delta k=k-K_{0}$ as $(\Delta k)^{-B}$, which scales with $\Delta K_{\text{min}}\sim L^{B}$, the model-dependent exponent $B$ related to the order of the EP [63, 66]. Such divergences in $P_{k}$ have also been recently shown to lead to $S\sim\log L$ entanglement scaling, albeit with a negative coefficient determined by $B$  [63, 66], as presented in Supp. Sect. I .
@@
Another key non-Hermitian phenomenon is the non-Hermitian skin effect (NHSE), but below we explain why the critical $\log L$ EE scaling still generically holds, as shown for the prototypical gapless non-Hermitian SSH [Fig. 1(c)] and other longer-ranged asymmetric hopping models I. In general, NHSE "skin" states can be modeled with exponentially decaying state profiles $\text{e}^{ikx}\text{e}^{-\kappa(k)x}$ for each wavenumber $k$, the inverse decay length $\kappa(k)$ determined through detailed boundary condition analysis [96, 99] [^4] This can involve various subtleties when there is more than one competing NHSE channel, as in our entanglement dip model discussed later.. As such, the periodic and open boundary condition (PBC and OBC) eigenstates are approximately related through a complex quasi-momentum deformation $k\rightarrow K(k)=k+i\kappa(k)$, such that $\psi^{R}_{k,\mu}\rightarrow\psi^{R}_{K,\mu}$. This path $K(k)$, $k\in[0,2\pi)$ is known as the generalized Brillouin zone (GBZ), and can be visualized as a closed loop $z=\text{e}^{iK(k)}$ that deviates from the unit circle $\text{e}^{ik}$ due to the NHSE, as illustrated in Fig. 1(c).
@@
Importantly, the modified entanglement entropy due to the NHSE can be directly computed through this GBZ deformation, since the eigen-equations for the left and right eigenstates depend on $K(k)$ according to  [^5] In Eq. 4, $\mathcal{H}$ and $E_{\mu}$ are regarded as functions of $K$, such that the conjugation operator only acts on the coefficients of $K$, and not $K$ itself.
@@
such that in the presence of skin state accumulation, $P_{k}\rightarrow\sum\limits_{\mu\in occ.}|\psi^{R}_{K,\mu}\rangle[|\psi^{L}_{K^{*},\mu}\rangle]^{\dagger}=P_{K}$. As such, replacing $P_{k}$ in Eq. LABEL:barP by $P_{K}$, we obtain the 2-point functions under the NHSE as
@@
where $\Delta K_{\text{min}}$ is determined  [^6] To satisfy OBCs, the discrete effective momenta on the GBZ are obtained by subdividing $2\pi$ into $L+1$ instead of $L$ intervals, and discarding the endpoints. by the closest approach of $K(2\pi n/(L+1))$, $n=1,...,L$ to the singularity. Since $\Delta K_{\text{min}}$ would still scale like $L^{-1}$ (brown curve) except in rare cases with sufficiently badly behaved $K(k)$  [^7] Even when the GBZ $K(k)=k+i\kappa(k)$ contains cusps, the singularities encountered by $P_{K}(k)$ at most acquire different branching numbers, and that can only modify the coefficient of the $\log L$ EE scaling., the key takeaway is that the NHSE itself cannot give rise to critical entanglement scaling other than $S\sim\log L$.
@@
In this work, a key insight is that qualitatively new entanglement scaling can however occur if the GBZ trajectory $K(k)$ also depends on $L$ i.e. is scale-dependent. As $L$ increases, $\Delta K_{\text{min}}$ may vary non-monotonically with $L$, reaching a minimum when the $K(k)$ cuts across a singularity $K_{0}$, as highlighted in Fig. 1(d) in red. The GBZ loop $\text{e}^{iK(2\pi n/(L+1))}$, $n=1,...,L$ expands with increasing $L$ (outward cyan arrows on loop of black dots), such that it intersects $K_{0}$ (dark purple) only at a special value of $L=L_{c}$. As such, we have rapid unconventional $\Delta K_{\text{min}}\propto|L-L_{c}|$ scaling around $L\approx L_{c}$ (kink on brown curve). Since $S\sim-\log\Delta K_{\text{min}}$, we thus also expect $\log L$-violating EE scaling.
@@
Condition (i) can be generically met by weakly coupling two finite 1D NHSE chains $\nu=\text{I}$, II with oppositely directed NHSE pumping to induce inter-chain tunneling [Fig. 2(a)]. As an important departure from usual literature, we label their basis $|x\rangle_{\nu},|k\rangle_{\nu}$ in opposite directions [Fig. 2(a)], such that both chains are described by the same Hamiltonian $H$. This labeling, which assigns the momentum $k$ based on the NHSE pumping, would turn out instrumental in our GBZ construction. Our system, with weak coupling $\delta$, takes the form
@@
where the last line describes only the bulk structure. Crucially, skin states grow exponentially with distance $L$ in each chain, concomitantly enhancing the effective inter-chain tunneling probability. At small $L$, we have two effectively uncoupled OBC chains (faint double-arrows in Fig. 2(a)). But at large $L$, the exponentially larger tunneling probabilities (dark blue double-arrows) effectively close the two antagonistic chains into a loop, forming a PBC-like configuration. This construction indeed produces a distinctively scale-dependent spectrum, as shown in Fig. 2(b) for a minimal illustrative model with coupled
@@
chains, where $\sigma_{\pm}=(\sigma_{x}\pm i\sigma_{y})/2$ are the Pauli matrices (see Fig. 4 and the supplement sect. I for other examples).
Figure 2: Scaling-dependent critical spectrum and GBZ from antagonistic NHSE. (a) Our construction [Eq. 7] involves weakly coupling two chains $H$ with opposite NHSE and momentum label $k$. Only for sufficiently long chains (Right) is the NHSE strong enough to effectively close the two chains into a loop with PBC-like dynamics. (b) Strongly $L$-dependent spectrum which undergoes a gapped to gapless exceptional critical transition at $L=47\approx L_{c}$ (cyan). (c) The corresponding $L$-dependent GBZ $z=\text{e}^{iK(k)}$ (colored loops), which pass through the critical point $K_{0}$ (where phase rigidity $r(z)=0$) when $L$ sweeps pass $L_{c}$. Parameters are $\delta=8.4\times 10^{-4}$ and $t_{L}=1.62$, $t_{R}=0.89$ in Eq. 8.
While no exact analytic solution exists for the GBZ of generic $\mathcal{H}$ [Eq. 7], below we outline a general procedure for deriving an approximate GBZ that nevertheless captures its scaling properties accurately:
@@
    Introduce the inter-chain coupling $\delta$ and solve the OBC eigenequation |$\mathcal{H}\ket{\Psi_{\text{GBZ}}}=E\ket{\Psi_{\text{GBZ}}}$ for both coupled chains using the ansatz |$\ket{\Psi_{\text{GBZ}}}=\sum_{i}c_{i}\sum\limits_{x=1}^{L}z_{i}^{x}\ket{\zeta(z_{i})}_{\text{I}}\oplus z_{i}^{L+1-x}\ket{\zeta(z_{i})}_{\text{II}}$, where $c_{i}$ are linear combination coefficients to be eliminated. The idea is to express the $\delta$ coupling terms entirely in terms of the OBC boundary contributions, both of which are what did not exist in the single-chain bulk solution in (2) above.
@@
    From the above, solve for the full GBZ solutions $z_{i}$ in terms of $L$, $\delta$ and $H$ model parameters, eliminating $E$ using the bulk dispersion from (1) if necessary. Due to the $z_{i}^{L+1-x}$ factor in our ansatz, which arises from the antagonistic momenta in Eq. 7, the GBZs $z_{i}$ typically satisfy a $L$-degree polynomial involving the model parameters, leading to generic $\sim 1/L$-th power scaling of $z_{i}$ with $\delta$.
This GBZ effectively encapsulates the non-perturbative effects of the coupling $\delta$ through the scale-dependent complex deformation $K(k)$, henceforth allowing properties of the full system $\mathcal{H}$ to be computed just through a single chain $H$. As detailed in Supplement sect. I and plotted in [Fig. 2(c)], the dominant scale-dependent GBZ for our minimal example $H_{\text{SSH}}$ [Eq. $8$The subdominant GBZ solutions $|z_{i}|<|z_{1}|$ can control the lower bound of the decaying eigenstates, as detailed in Supplement Sect. I scales like
@@
$\alpha_{0}$ a model-dependent constant. Eq. 9 accurately predicts the spectrum of $\mathcal{H}(k)$ upon substituting $k\rightarrow-i\log z_{1}$ [Fig. 2(b)]. Correctly associating the antagonistic directions with opposite momenta [Eq. 7] was crucial in deriving this scaling-dependent GBZ and spectrum, which cannot be analytically derived as a function of $L$ using conventional approaches [102, 105, 145, 118, 112, 146].
Figure 3: Entanglement dip from SIEC. (a) Spikes in the $\bar{P}$ eigenvalues $p$ occur near $L\approx L_{c}$ [Eq. 10] for different inter-chain couplings $\delta=0.83\times 10^{-3},1.6843\times 10^{-3}$ and $3.03\times 10^{-3}$ (green, black and purple respectively) of Eq. 8, with hopping asymmetries $t_{L},t_{R}=1.2e^{\pm 0.3}$. (b) Corresponding entanglement dips at these $L\approx L_{c}$, with characteristic $S\sim(L_{c}-L)^{-1/2}$ behavior for $L<L_{c}$. (c) Depth of entanglement dips in the parameter space of $t_{L}/t_{R},\delta$ and the system sizes $L$ realizing them, with large dips of $S_{\text{min}}\leq-3$ along narrow bands separating regions of different $L$.
SIEC and entanglement dip.– Our $L$-dependent GBZ can give rise to a new phenomenon dubbed “scaling-induced exceptional criticality” (SIEC) if the GBZ loop $\text{e}^{iK(k)}$ sweeps through an EP (where $r(K)=0$, dark purple in Fig. 2(c)) when $L$ is varied across a special value $L_{c}$. For our illustrative $H_{\text{SSH}}$ and related models I, substituting Eq. 9 into $H_{\text{SSH}}(K(k))$ and demanding that it reduces to the Jordan form $\sigma_{-}$ at the EP yields
@@
[…]
@@
such that I $H_{\text{SSH}}(K)\approx(t_{L}-t_{R}^{-1})\sigma_{+}-it_{R}\Delta_{K}\sigma_{-}$ for $L\approx L_{c}$, with the expected square root cusp in its eigenenergies ±$\pm\mathcal{E}_{\text{SSH}}(K)=\pm\sqrt{i(1-t_{L}t_{R})\Delta_{K}}$. In particular, its occupied band projector, which assumes the form $P(K)=(\mathbb{I}-H(K)/\mathcal{E}(K))/2$ for 2-component models, has divergent off-diagonal term $-\sqrt{i(t_{L}-t_{R}^{-1})/4t_{R}\Delta_{K}}\,\sigma_{+}$. Hence, with $k=\pi$, we obtain the unique scaling divergence

### v2

@@ Abstract
Introduction.– Entanglement scaling behavior provides an established diagnostic for critical phase transitions. Deeply rooted in boundary conformal field theory [1], it universally associates logarithmic entanglement entropy (EE) scaling $S\sim\log L$ with 1D criticality [2, 3]. This seminal scaling property has also seen generalizations to higher-dimensional and topological systems, where midgap modes and nodal structures contribute distinctive signatures to the entanglement scaling, some with measurable prospects.
Of late, established results in critical entanglement scaling have been challenged in non-Hermitian contexts. In so-called exceptional points (EPs) [4], geometric defectiveness blurs the distinction between occupied and unoccupied bands and probability non-conservation leads to fermionic occupancies effectively greater than one, giving rise to enigmatically negative free-fermion EE through the 2-point function [6, 5]. This is further exacerbated by the non-Hermitian skin effect (NHSE), which produces macroscopically many highly defective eigenstates indexed by complex momenta. Yet, despite these complications, all known non-Hermitian free-fermion critical systems still exhibit the celebrated $S\sim\log L$ scaling.
@@
where $|k\rangle$ is the quasi-momentum basis state and $|\psi_{k,\mu}^{R}\rangle,\langle\psi^{L}_{k,\mu}|$ are respectively the right and left eigenbands [7] of $\mathcal{H}$ with band index $\mu$. We stipulate that fermions occupy an arbitrary set of bands $\mu\in occ.$, such that state occupancy is described by the occupied band projector $P=\sum_{k,\mu\in occ.}|\psi_{k,\mu}^{R}\rangle\langle\psi^{L}_{k,\mu}|=\sum_{k}P_{k}|k\rangle\langle k|$, where $P_{k}=\sum_{\mu\in occ.}\psi_{k,\mu}^{R}[\psi^{L}_{k,\mu}]^{*}$. For an entanglement cut where the region $[x_{L},x_{R}]$ is to be truncated, the biorthogonal free-fermion EE is given by [3]
@@
Based on Eq. (LABEL:barP), we first review why the EE scales like $\log L$ in Hermitian critical points, as illustrated in Fig. 1a. Consider a critical point $K_{0}$ (red) where the gap closes but the eigenspace remains full-rank i.e. non-defective. Due to the possible mixing of the occupied and unoccupied eigenstates at gap closure, the occupied band projector $P_{k}$ generically become discontinuous i.e. singular at the critical point. However, due to the discrete momentum spacing (black dots) which scales like $2\pi/L$ (cyan arrows) in a finite lattice, the momentum sum in Eq. (LABEL:barP) only samples $P_{k}$ down to a distance $\Delta k=k-K_{0}$ bounded below by $\Delta K_{\text{min}}\sim\pi/L$ (brown curve) from the critical point i.e. $\bar{P}_{xx^{\prime}}\sim\int_{\pi/L}\text{e}^{i\Delta k(x^{\prime}-x)}P_{\Delta k}\,d(\Delta k)$ with a UV cutoff $\pi/L$. Since analytic singularities lead to power-law decay in their Fourier coefficients, we expect $P_{xx^{\prime}}$, which are also the 2-point functions, to become long-ranged and cause substantial entanglement across the cut. Mathematically, this entanglement can be shown to diverge logarithmically with the inverse of the UV cutoff, thereby yielding $S\sim\log L$ entanglement scaling.
@@
Such $\log L$ entanglement scaling persists even if the critical point is geometrically defective (with completely non-orthogonal eigenstates), as is possible in non-Hermitian systems, as illustrated in Fig. 1b. For a particular band, the overlap between its left and right eigenstates $\langle\psi^{L}_{k}|$, $|\psi^{R}_{k}\rangle$ can be quantified through the phase rigidity [8] $r(k)=\frac{\langle\psi^{L}_{k}|\psi^{R}_{k}\rangle}{\sqrt{\langle\psi^{L}_{k}|\psi^{L}_{k}\rangle\langle\psi^{R}_{k}|\psi^{R}_{k}\rangle}},$ which is always unity in Hermitian systems, but ranging between $0\leq r(k)\leq 1$ in non-Hermitian systems. In particular, at an exceptional critical point (EP) $K_{0}$ where $r(K_{0})=0$ [9] (dark purple), the occupied and unoccupied eigenstates coalesce and the occupied band projector $P_{k}$ is not just discontinuous, but in fact divergent. Specifically, one or more matrix elements of $P_{k}$ can be shown to diverge with $\Delta k=k-K_{0}$ as $(\Delta k)^{-B}$, which scales with $\Delta K_{\text{min}}\sim L^{B}$, the model-dependent exponent $B$ related to the order of the EP [5, 10]. Such divergences in $P_{k}$ have also been recently shown to lead to $S\sim\log L$ entanglement scaling, albeit with a negative coefficient determined by $B$  [5, 10], as presented in the Supplementary material Section II.
@@
Another key non-Hermitian phenomenon is the non-Hermitian skin effect (NHSE), but below we explain why the critical $\log L$ EE scaling still generically holds, as shown for the prototypical gapless non-Hermitian SSH [Fig. 1c] and other longer-ranged asymmetric hopping models. In general, NHSE "skin" states can be modeled with exponentially decaying state profiles $\text{e}^{ikx}\text{e}^{-\kappa(k)x}$ for each wavenumber $k$, the inverse decay length $\kappa(k)$ determined through detailed boundary condition analysis [6, 11]. As such, the periodic and open boundary condition (PBC and OBC) eigenstates are approximately related through a complex quasi-momentum deformation $k\rightarrow K(k)=k+i\kappa(k)$, such that $\psi^{R}_{k,\mu}\rightarrow\psi^{R}_{K,\mu}$. This path $K(k)$, $k\in[0,2\pi)$ is known as the generalized Brillouin zone (GBZ), and can be visualized as a closed loop $z=\text{e}^{iK(k)}$ that deviates from the unit circle $\text{e}^{ik}$ due to the NHSE, as illustrated in Fig. 1c.
@@
Importantly, the modified entanglement entropy due to the NHSE can be directly computed through this GBZ deformation, since the eigen-equations for the left and right eigenstates depend on $K(k)$ according to
@@
such that in the presence of skin state accumulation, $P_{k}\rightarrow\sum\limits_{\mu\in occ.}|\psi^{R}_{K,\mu}\rangle[|\psi^{L}_{K^{*},\mu}\rangle]^{\dagger}=P_{K}$. As such, replacing $P_{k}$ in Eq. (LABEL:barP) by $P_{K}$, we obtain the 2-point functions under the NHSE as
@@
where $\Delta K_{\text{min}}$ is determined by the closest approach of $K(2\pi n/(L+1))$, $n=1,...,L$ to the singularity. Since $\Delta K_{\text{min}}$ would still scale like $L^{-1}$ (brown curve) except in rare cases with sufficiently badly behaved $K(k)$, the key takeaway is that the NHSE itself cannot give rise to critical entanglement scaling other than $S\sim\log L$.
@@
In this work, a key insight is that qualitatively new entanglement scaling can however occur if the GBZ trajectory $K(k)$ also depends on $L$ i.e. is scale-dependent. As $L$ increases, $\Delta K_{\text{min}}$ may vary non-monotonically with $L$, reaching a minimum when the $K(k)$ cuts across a singularity $K_{0}$, as highlighted in Fig. 1d in red. The GBZ loop $\text{e}^{iK(2\pi n/(L+1))}$, $n=1,...,L$ expands with increasing $L$ (outward cyan arrows on loop of black dots), such that it intersects $K_{0}$ (dark purple) only at a special value of $L=L_{c}$. As such, we have rapid unconventional $\Delta K_{\text{min}}\propto|L-L_{c}|$ scaling around $L\approx L_{c}$ (kink on brown curve). Since $S\sim-\log\Delta K_{\text{min}}$, we thus also expect $\log L$-violating EE scaling.
@@
Condition (i) can be generically met by weakly coupling two finite 1D NHSE chains $\nu=\text{I}$, II with oppositely directed NHSE pumping to induce inter-chain tunneling [Fig. 2a]. As an important departure from usual literature, we label their basis $|x\rangle_{\nu},|k\rangle_{\nu}$ in opposite directions [Fig. 2a], such that both chains are described by the same Hamiltonian $H$. This labeling, which assigns the momentum $k$ based on the NHSE pumping, would turn out instrumental in our GBZ construction. Our system, with weak coupling $\delta$, takes the form
@@
where the last line describes only the bulk structure. Crucially, skin states grow exponentially with distance $L$ in each chain, concomitantly enhancing the effective inter-chain tunneling probability. At small $L$, we have two effectively uncoupled OBC chains (faint double-arrows in Fig. 2a). But at large $L$, the exponentially larger tunneling probabilities (dark blue double-arrows) effectively close the two antagonistic chains into a loop, forming a PBC-like configuration. This construction indeed produces a distinctively scale-dependent spectrum, as shown in Fig. 2b for a minimal illustrative model with coupled
@@
chains, where $\sigma_{\pm}=(\sigma_{x}\pm i\sigma_{y})/2$ are the Pauli matrices (see Fig. S13 and the Supplementary material for other examples).
Figure 2: Scaling-dependent critical spectrum and GBZ from antagonistic NHSE. (a) Our construction [ Eq. (7)] involves weakly coupling two chains $H$ with opposite NHSE and momentum label $k$. Only for sufficiently long chains (Right) is the NHSE strong enough to effectively close the two chains into a loop with PBC-like dynamics. (b) Strongly $L$-dependent spectrum which undergoes a gapped to gapless exceptional critical transition at $L=47\approx L_{c}$ (cyan). (c) The corresponding $L$-dependent GBZ $z=\text{e}^{iK(k)}$ (colored loops), which pass through the critical point $K_{0}$ (where phase rigidity $r(z)=0$) when $L$ sweeps pass $L_{c}$. Parameters are $\delta=8.4\times 10^{-4}$ and $t_{L}=1.62$, $t_{R}=0.89$ in Eq. (8).
While no exact analytic solution exists for the GBZ of generic $\mathcal{H}$ [ Eq. (7)], below we outline a general procedure for deriving an approximate GBZ that nevertheless captures its scaling properties accurately:
@@
    Introduce the inter-chain coupling $\delta$ and solve the OBC eigenequation |$\mathcal{H}\ket{\Psi_{\text{GBZ}}}=E\ket{\Psi_{\text{GBZ}}}$ for both coupled chains using the ansatz |$\ket{\Psi_{\text{GBZ}}}=\sum_{i}c_{i}\sum\limits_{x=1}^{L}z_{i}^{x}\ket{\zeta(z_{i})}_{\text{I}}\oplus z_{i}^{L+1-x}\ket{\zeta(z_{i})}_{\text{II}}$, where $c_{i}$ are linear combination coefficients to be eliminated. The idea is to express the $\delta$ coupling terms entirely in terms of the OBC boundary contributions, both of which are what did not exist in the single-chain bulk solution in step 2 above. Although $\delta$ is very small, the coupling effect is non-perturbative due to the strong boundary sensitivity of non-Hermitian systems with the skin effect.
@@
    From the above, solve for the full GBZ solutions $z_{i}$ in terms of $L$, $\delta$ and $H$ model parameters, eliminating $E$ using the bulk dispersion from (1) if necessary. Due to the $z_{i}^{L+1-x}$ factor in our ansatz, which arises from the antagonistic momenta in Eq. (7), the GBZs $z_{i}$ typically satisfy a $L$-degree polynomial involving the model parameters, leading to generic $\sim 1/L$-th power scaling of $z_{i}$ with $\delta$.
This GBZ effectively encapsulates the non-perturbative effects of the coupling $\delta$ through the scale-dependent complex deformation $K(k)$, henceforth allowing properties of the full system $\mathcal{H}$ to be computed just through a single chain $H$. As detailed in the Supplementary material and plotted in [Fig. 2c], the dominant scale-dependent GBZ for our minimal example $H_{\text{SSH}}$ [ Eq. (8)] scales like
@@
$\alpha_{0}$ a model-dependent constant. Eq. (9) accurately predicts the spectrum of $\mathcal{H}(k)$ upon substituting $k\rightarrow-i\log z_{1}$ [Fig. 2b]. Correctly associating the antagonistic directions with opposite momenta [ Eq. (7)] was crucial in deriving this scaling-dependent GBZ and spectrum, which cannot be analytically derived as a function of $L$ using conventional approaches [15].
Figure 3: Entanglement dip from SIEC. (a) Spikes in the $\bar{P}$ eigenvalues $p$ occur near $L\approx L_{c}$ [ Eq. (10)] for different inter-chain couplings $\delta=0.83\times 10^{-3},1.6843\times 10^{-3}$ and $3.03\times 10^{-3}$ (green, black and purple respectively) of Eq. (8), with hopping asymmetries $t_{L},t_{R}=1.2e^{\pm 0.3}$. (b) Corresponding entanglement dips at these $L\approx L_{c}$, with characteristic $S\sim(L_{c}-L)^{-1/2}$ behavior for $L<L_{c}$. (c) Depth of entanglement dips in the parameter space of $t_{L}/t_{R},\delta$ and the system sizes $L$ realizing them, with large dips of $S_{\text{min}}\leq-3$ along narrow bands separating regions of different $L$.
SIEC and entanglement dip.– Our $L$-dependent GBZ can give rise to a new phenomenon dubbed “scaling-induced exceptional criticality” (SIEC) if the GBZ loop $\text{e}^{iK(k)}$ sweeps through an EP (where $r(K)=0$, dark purple in Fig. 2c) when $L$ is varied across a special value $L_{c}$. For our illustrative $H_{\text{SSH}}$ and related models, substituting Eq. (9) into $H_{\text{SSH}}(K(k))$ and demanding that it reduces to the Jordan form $\sigma_{-}$ at the EP yields
@@
which accurately predicts when the exceptional transition occurs, corroborated by numerical results in Fig. 2b. As $L_{c}$ is typically non-integer, actual lattice $\mathcal{H}$ models with integer $L$ experience dramatic SIEC divergences when $L\rightarrow L_{c}$, as controlled by
@@
such that $H_{\text{SSH}}(K)\approx(t_{L}-t_{R}^{-1})\sigma_{+}-it_{R}\Delta_{K}\sigma_{-}$ for $L\approx L_{c}$, with the expected square root cusp in its eigenenergies ±$\pm\mathcal{E}_{\text{SSH}}(K)=\pm\sqrt{i(1-t_{L}t_{R})\Delta_{K}}$. In particular, its occupied band projector, which assumes the form $P(K)=(\mathbb{I}-H(K)/\mathcal{E}(K))/2$ for 2-component models, has divergent off-diagonal term $-\sqrt{i(t_{L}-t_{R}^{-1})/4t_{R}\Delta_{K}}\,\sigma_{+}$. Hence, with $k=\pi$, we obtain the unique scaling divergence